# Improving Factuality for Dialogue Response Generation via Graph-Based Knowledge Augmentation

### v1

@@ Abstract
Large Language Models (LLMs) succeed in many natural language processing tasks. However, their tendency to hallucinate — generate plausible but inconsistent or factually incorrect text — can cause problems in certain tasks, including response generation in dialogue. To mitigate this issue, knowledge-augmented methods have shown promise in reducing hallucinations. Here, we introduce a novel framework designed to enhance the factuality of dialogue response generation, as well as an approach to evaluate dialogue factual accuracy. Our framework combines a knowledge triple retriever, a dialogue rewrite, and knowledge-enhanced response generation to produce more accurate and grounded dialogue responses. To further evaluate generated responses, we propose a revised fact score that addresses the limitations of existing fact-score methods in dialogue settings, providing a more reliable assessment of factual consistency. We evaluate our methods using different baselines on the OpendialKG and HybriDialogue datasets. Our methods significantly improve factuality compared to other graph knowledge-augmentation baselines, including the state-of-the-art G-retriever. The code will be released on GitHub  [^1] https://github.com/XiangyanChen/Triton.
Improving Factuality for Dialogue Response Generation via Graph-Based Knowledge Augmentation
Xiangyan Chen1  Yujian Gan1  Matthew Purver1,2 1Queen Mary University of London, UK 2Institut Jožef Stefan, Slovenia {xiangyan.chen, y.gan, m.purver}@qmul.ac.uk
@@ 1 Introduction
Figure 1: Examples comparing an LLM’s dialogue response generation when utilising retrieved triples and a rewritten dialogue versus generating responses without external knowledge or dialogue rewriting. The knowledge triple retriever module assists in selecting the triples most relevant to the query.
@@
Large Language Models (LLMs) have been shown to perform powerfully on Natural Language Processing (NLP) tasks. Despite their general superiority, LLMs will generate some plausible but fact-inconsistent text, namely hallucination. Flawed pre-training data, model bias and randomness in inference are the factors contributing to hallucinations Zhang et al. (2023); Huang et al. (2023). In dialogue response generation, generating incorrect responses will mislead people and have a negative impact on society.
@@
A number of methods have been proposed to enhance the factuality of language models. Among them, knowledge-augmented inference methods aim to do this by adding knowledge triples into the prompt. This has been effective in tasks such as question-answering Baek et al. (2023); Sen et al. (2023); Wu et al. (2023).
@@
However, these works Baek et al. (2023); Wu et al. (2023) tend to use simple sentence embedders to encode knowledge and determine the match between a query and knowledge based on similarity, rather than employing a specialised model designed to assess the relevance between the query and knowledge. He et al. (2024) proposed a graph-based method called G-retriever, which employs the Prize-Collecting Steiner Tree to help select knowledge triples. However, applying these methods to dialogue response generation is challenging due to their lack of consideration for dialogue context. Dialogues frequently involve intricate coreference structures, which complicate entity linking and hinder the LLMs’ comprehension, ultimately compromising the factual consistency and quality of the generated responses.
To address above limitations, we propose a novel framework to improve dialogue factuality that consists of three key components: a knowledge triple retriever, a dialogue rewriting and knowledge-enhanced dialogue response generation. The knowledge triple retriever is fine-tuned on our well-collected samples, which helps select the valuable triples for the dialogue. The dialogue rewrite is based on Chain of Thought (CoT), which resolves the dialogue coreference. The knowledge-enhanced dialogue generation is proposed to improve the dialogue factuality, based on two kinds of approaches: prompt-based and graph-based. Figure 1 illustrates how the knowledge-augmented generation works. The query asks about what movies star “Diane Lane”, and the initial response generated by the LLM contains factual inaccuracies. After applying a dialogue rewrite module that resolves coreference, a triple retriever helps select the most relevant N triples, including (“Nights in Rodanthe”, “cast member”, “Diane Lane”). With these corrected triples, the model generates an accurate response.
@@
To evaluate factuality for responses, we primarily rely on the fact score Min et al. (2023). However, the original design of the fact score does not consider dialogue context and situations where the knowledge source is unavailable, making it difficult to evaluate dialogue responses. To address this limitation, we adapt the fact score to the dialogue settings and assess model-human agreement. Annotation results show that our adapted fact score achieves substantial model-human agreement. Additionally, we propose Not Enough Information Proportion (NEIP) to evaluate dialogue factuality comprehensively. It is a metric that quantifies the proportion of atomic facts in a response that cannot be verified, such as opinions or hallucinated content that can not find any evidence to verify. We utilize the F1 score as a supplementary metric to assess factual accuracy and evaluate response quality through BLEU Reiter (2018), ROUGE-L Lin (2004), and Perplexity (PPL) Jelinek et al. (1977). Additionally, we perform human evaluations to further assess response quality.
@@
We compare our proposed method not only with standalone LLMs but also with knowledge-enhanced approaches such as KAPING, the BM25 algorithm, and the State-of-the-Art (SOTA) G-retriever He et al. (2024). Experimental results show that our framework consistently outperforms both existing knowledge-enhanced methods and the current best-performing G-retriever.
@@
    We propose a novel framework designed to enhance the factual accuracy of dialogue response generation. This framework incorporates three key components: a knowledge triple retriever that selects valuable triples, a dialogue rewrite that resolves coreference, and a knowledge-enhanced mechanism that effectively boosts overall factuality.
@@
    We adapt the fact score, originally designed for biography generation, to evaluate the factuality of dialogue systems, ensuring its validity within our task. The evaluation is fine-grained, enabling a more reliable assessment of factual consistency in dialogue systems.
@@
    We validate our methods against various baselines using two public dialogue datasets. Experimental results demonstrate that our approach significantly outperforms existing methods in factuality while keeping good quality at the generated response.
@@ 2.1 LLM Hallucinations and Mitigation Methods
LLMs are trending in NLP, whose architecture can mainly be classified as encoder-only Devlin (2018), decoder-only Brown (2020); Dubey et al. (2024), and encoder-decoder Chung et al. (2024). However, hallucination exists in these LLMs widely, which is the phenomenon that language models generate coherent but fact-inconsistent text. Many reasons contribute to hallucination. Huang et al. (2023) declared one potential reason is that the data in the pre-training of LLMs is incomplete, incorrect or outdated. Other reasons for hallucinations are the bias of LLMs and randomness in inference. Zhang et al. (2023) categorise hallucinations as input-conflicting, context-conflicting and fact-conflicting, with recent focus on the latter.
Several knowledge-based methods have been proposed to mitigate fact-conflicting hallucinations Agrawal et al. (2023), which generally fall into three categories: knowledge-aware inference, knowledge-aware training, and knowledge-aware validation. For instance, Baek et al. (2023) introduced KAPING, a prompt-based framework for question answering that retrieves knowledge triples based on embedding similarity. Building upon retrieval techniques, Sen et al. (2023) combined Knowledge Graph (KG) retrieval with language model reasoning to enhance performance on complex questions. Similarly, CRAG Yan et al. (2024) employs an evaluator to assess generation quality and, if necessary, refines outputs via web search, showing significant improvements across various generation tasks. Another notable approach is Self-RAG Asai et al. (2023), which retrieves relevant knowledge, generates an initial answer, and then refines it through self-critique, achieving strong results in multiple question answering benchmarks. In a related development, He et al. (2024) proposed G-Retriever, which integrates a graph-based encoder with the Prize-Collecting Steiner Tree algorithm to encode retrieved graph knowledge effectively, standing out among graph-enhanced knowledge retrieval methods.
@@ 2.2 Evaluation of Factuality
Dialogue response generation differs from question answering in that it must not only take the dialogue context into account when generating responses, but also must be evaluated based on the factuality and quality of the responses, rather than just the accuracy of answers. Previous dialogue factuality evaluation mainly focuses on humans Ni et al. (2023); Li et al. (2022); Yu et al. (2022), and it is inefficient.
@@
Several question-answering datasets provide factual references in the form of entities; reference-based evaluation can then be based directly on entity matching. Other NLP datasets offer text references, so an alternative reference-based method involves matching the extracted entities between the generated text and these references Nan et al. (2021). Instead, dialogue datasets provide reference responses. To evaluate the factual accuracy of the generated text, we can therefore compare the entities extracted from the generated response with those in the reference response, measured as F1 score. However, relying solely on this can be inadequate, as the reference may not always provide explicit answers. It is crucial to consider additional metrics to ensure a comprehensive evaluation.
Reference-free methods can be categorised into uncertainty estimation Farquhar et al. (2024) and external knowledge-based approaches, with the former often limited by its reliance on the generation model’s own confidence. The fact score Min et al. (2023), which falls into the latter category, is a metric designed to measure fact consistency in long-form text. The process begins with an LLM breaking down the text into fine-grained sentences called atomic facts. These atomic facts are then verified using both the LLM and external knowledge sources (Min et al. (2023) use Wikipedia titles). The fact score is calculated based on the precision.
However, the fact score is not directly suitable for our work for two key reasons: First, it evaluates only the generated text without taking the dialogue history into account, which is crucial when assessing the factuality of dialogue responses. Second, it does not account for the possibility that generated responses may lack external knowledge sources for verification.
Figure 2: The workflow of dialogue response generation and factuality evaluation. The left part of the figure is our framework. It starts with rewriting the dialogue, and then the knowledge triples will be retrieved from KG. The triple retriever helps select valuable triples. The colour-coded arrows marked as Serial Number 6 represent different knowledge-enhanced methods: the orange arrow indicates the graph-based approach, while the blue arrow corresponds to the prompt-based approach. The right part is factuality evaluation, in which the response will be broken down into individual and verifiable facts, checked using both the LLM and external knowledge sources.
@@ 3 Methodology
The diagram in Figure 2 illustrates the workflow of dialogue generation (left side) and factuality evaluation (right side).
Our proposed framework consists of three key components: (1) A well-designed knowledge triple retriever and see Section 3.2 for details, (2) A CoT-based dialogue rewriter, described in Section 3.3, and (3) the knowledge-enhanced dialogue response generation (Section 3.2).
For factuality evaluation (Section 3.5), we adopt the fact score as a primary metric. However, as the original fact score does not fully capture certain aspects of dialogue evaluation, we introduce modifications to better align with our objectives.
@@ 3.1 Task Formulation
Our task is to generate a response $s$ given a dialogue context $c$, query $x$ and set of triples $f=\{(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),...,(h_{n},r_{n},t_{n})\}$ from an encyclopedia-based KG, where $h$ and $t$ represent the head and tail entities, and $r$ denotes the relation between them.
3.2 Knowledge Retriever
The KG is a collection of triples, and these triples are the form of head entity, relation, and tail entity. In encyclopedia-based KGs, like Wikidata Vrandečić and Krötzsch (2014), a triple usually indicates a fact. We aim to retrieve triples from the KG for the query in dialogue and incorporate triples into prompts as a supplement.
We first extract entities from the query using the entity linking approach, and then, we match all entities with the triples from the KG when the entity equals the head or tail entity. Since our entities are sourced from Wikidata, and ReFinED Ayoola et al. (2022) is an entity linking method based on corresponding Wikidata entities, we adopt it for our retrieval process.
Knowledge retriever gathers all triples associated with the entities extracted from the query, but not all of these are relevant to the query. Irrelevant triples introduce textual noise, which has a negative impact on text generation. Thus, it is crucial to select the most relevant triples for the query. Traditional methods rank triples based on word frequency using BM25 Robertson et al. (2009), which primarily captures word overlap. Alternatively, sentence embedding models like MPNet Song et al. (2020), as employed by KAPING Baek et al. (2023), encode semantic information but in a relatively straightforward manner.
To retrieve delicate knowledge, we collect query-triple pairs to fine-tune LLM as a knowledge matcher.
Data Collection
Manual data annotation is both time-consuming and labour-intensive. Given GPT’s strong performance in NLP tasks, we adopt it as an annotation tool. Our approach extracts query-triple pairs from the training dataset and prompts GPT-4o to determine their relevance. The model classifies each pair as either relevant or irrelevant. Based on our manual assessment, GPT-4o demonstrates high accuracy in determining query-triple relevance, described in Appendix F.
Fine-Tuning
After collecting these data, we fine-tune an LLM as the binary classifier. In our work, we select the Llama3 8B model Dubey et al. (2024). We combined LoRA Hu et al. (2021) with fine-tuning, which is an efficient fine-tuning method that adopts low-rank modules into transformer layers. Cross-entropy loss is used to optimize the model during fine-tuning, and it is formulated as follows:
$\mathcal{L}=-\sum_{i=1}^{N}y_{i}\log(p_{i})$ (1)
where $y_{i}$ denotes the true label, $p_{i}$ is defined as the predicted probability and $N$ is the number of label classifications.
With the fine-tuned LLM $\mathcal{M}_{t}$, the query $x$ and the triple $f_{i}$, the probability of relevance can be defined as:
$P(y=\text{relevant}|x,f_{i})=\mathcal{M}_{\text{t}}(x,f_{i})$ (2)
Then, the selected triples are formulated as:
$f_{\text{sorted}}=\text{sort}_{\text{desc}}(P,f,x)$ (3)
where $\text{sort}_{\text{desc}}$ denotes the sort function descending by $P$, and $P$ is the relevance probability distribution.
3.3 Dialogue Rewrite
Unlike in question answering tasks, dialogue response generation involves coreferences that pose challenges for both entity linking and response generation. An intuitive approach to improve entity linking is to expand the scope for extracting entities from the dialogue context. However, this approach may also introduce irrelevant entities, negatively impacting performance.
To resolve coreference and improve the performance of entity linking and dialogue response generation, we propose a CoT-based coreference resolution.
Data Collection
We prompt GPT-4o with dialogue to resolve the conversation in a zero-shot CoT setting, as described by Kojima et al. (2022). Step-by-step reasoning in LLMs enhances the generation of accurate results and aids in fine-tuning smaller language models.
Fine-Tuning
We fine-tune the Llama3 8B model using LoRA on collected coreference resolution samples. The reason for selecting LoRA for this task is that it will not destroy the LLMs reasoning capability obtained from pre-training, making it perform better on this task. Cross-entropy loss is also selected for this task.
After fine-tuning, we prompt LLM to resolve coreference, and the dialogue is rewritten. The rewrite and corresponding sorted triples are formulated as follows:
$c^{\prime},x^{\prime}=\mathcal{M}_{\text{f}}(p_{\text{rewrite}},c,x)$ (4)
$f_{\text{sorted}}^{{}^{\prime}}=\text{sort}_{\text{desc}}(P,f,x^{\prime})$ (5)
where $p_{\text{rewrite}}$, described in Table 6 of Appendix A, is the prompt for rewriting dialogue. $\mathcal{M}_{\text{f}}$ is the LLM for dialogue rewriting.
3.4 Knowledge-Based Dialogue Response Generation
To mitigate hallucinations of LLM and have a factual response, we offer two patterns of generating dialogue responses given selected knowledge triples: prompt-based and graph-based methods.
Prompt Based Generation
Given dialogue context $c$, query $x$, rewritten dialogue $c^{\prime}$ and $x^{\prime}$, as well as selected triples $f_{sorted}$ and $f_{sorted}^{{}^{\prime}}$, the generation of knowledge-augmented dialogue responses and their counterparts with the rewritten dialogue are formulated as follows:
$s=\mathcal{M}_{\text{g}}(p_{\text{gen}},c,x,f_{\text{sorted}})$ (6)
$s^{{}^{\prime}}=\mathcal{M}_{\text{g}}(p_{\text{gen}},c^{\prime},x^{\prime},f_{\text{sorted}}^{{}^{\prime}})$ (7)
where $p_{\text{gen}}$ denotes the prompt for generating the response. $\mathcal{M}_{\text{g}}$ is the LLM for response generation. Table 6 in Appendix A describes the details of the prompt template.
Graph-Based Generation

### v2

# Improving Factuality for Dialogue Response Generation via Graph-Based Augmentation

@@ Abstract
Large Language Models (LLMs) succeed in many natural language processing tasks. However, their tendency to hallucinate — generate plausible but inconsistent or factually incorrect text — can cause significant problems in certain tasks, including response generation in dialogue. To mitigate this issue, we propose two novel graph knowledge-augmented frameworks, Dialogue Response Generation via Textualised Graphs (TG-DRG) and Graph-Aware Dialogue Response Generation (GA-DRG), which combine reasoning-guided dialogue reformulation, dialogue sense knowledge selection, and graph-enhanced response generation to improve the factuality of dialogue responses. To evaluate the factuality of generated responses, we propose a dialogue fact score that addresses the limitations of existing fact-score methods in dialogue settings, providing a more reliable assessment of factual consistency. We evaluate our methods using different baselines on the OpendialKG and HybriDialogue datasets. Our methods noticeably improve factuality compared to other graph knowledge-augmentation baselines, including the state-of-the-art G-retriever, achieving improvements of 3.47% on OpendialKG and 3.12% on HybriDialogue in terms of dialogue fact score. The code will be released on GitHub [^1] https://github.com/XiangyanChen/DRG.
@@ 1 Introduction
Large Language Models (LLMs) have been shown to perform powerfully on Natural Language Processing (NLP) tasks. Despite their general superiority, LLMs will generate some plausible but fact-inconsistent text, namely hallucination. Flawed pre-training data, model bias and randomness in inference are the factors contributing to hallucinations (Zhang et al. 2023; Huang et al. 2023). In dialogue response generation, generating incorrect responses will mislead people and have a negative impact on society.
@@
A number of approaches have been proposed to enhance the factuality of LLMs. Among them, graph knowledge-augmented methods aim to improve factual accuracy by incorporating knowledge graphs. This has been effective in tasks such as Question Answering (QA) (Baek, Aji, and Saffari 2023; Wu et al. 2023; He et al. 2024).
Figure 1: An example illustrates how our frameworks TG-DRG and GA-DRG, and dialogue fact score work.
@@
Recently, Baek, Aji, and Saffari (2023) utilised sentence embedders to encode knowledge and determine the match between a query and knowledge based on similarity. He et al. (2024) proposed a graph-based method called G-retriever, which employs the Prize-Collecting Steiner Tree (PCST) to help construct a subgraph. However, applying these methods to dialogue response generation is challenging due to their lack of consideration for dialogue context. Dialogues frequently involve intricate coreference structures, which hinder the LLMs’ comprehension, ultimately compromising the factual consistency and quality of the generated responses. In addition, evaluating factuality for dialogue responses is a challenge. F1-score relying on entities overlapping has limitations on open-domain tasks. More advanced FactScore (Min et al. 2023) evaluates facuality by splitting long-text into smaller atomic facts, but it does not consider dialogue context and situations where the knowledge source is unavailable, making it difficult to evaluate dialogue responses.
To overcome the limitation on generating factually correct responses, we propose two frameworks: (1) Dialogue Response Generation via Textualised Graphs (TG-DRG), where generation requires no training; and (2) Graph Aware Dialogue Response Generation (GA-DRG), which explicitly models node relationships, designed to improve the factual accuracy of dialogue generation. Both frameworks are built around reasoning-guided dialogue reformulation and dialogue sense knowledge selection. The latter helps select the most valuable knowledge based on the dialogue. We use an example shown in Figure 1 to illustrate how our frameworks work. The response to the dialogue is incorrect due to the pronouns, e.g. “her”. After applying Chain of Thought (CoT), the pronoun is solved, benefiting further knowledge-augmented dialogue response generation.
@@
To address limitations in evaluating factual accuracy, we propose the Dialogue Fact Score, as illustrated in Figure 1. In this case, the FactScore can not work without a dialogue context. Our metric accounts for both the dialogue context and scenarios in which the knowledge source is partially or entirely unavailable. Additionally, we propose Not Enough Information Proportion (NEIP) to evaluate dialogue factuality comprehensively. It is a metric that quantifies the proportion of atomic facts in a response that cannot be verified, such as opinions or hallucinated content that can not find any evidence to verify. We assess model-human agreement of dialogue fact score. Annotation results show that our dialogue fact score achieves substantial model-human agreement.
@@
We compare our proposed method not only with standalone LLMs but also with knowledge-enhanced approaches such as KAPING (Baek, Aji, and Saffari 2023), the BM25 algorithm, and the State-of-the-Art (SOTA) G-retriever (He et al. 2024). Experimental results show that our framework consistently outperforms both existing knowledge-enhanced methods and the current best-performing G-retriever.
@@
    We propose the dialogue fact score to evaluate the factuality of dialogue systems automatically. The evaluation is fine-grained, enabling a more reliable assessment of factual consistency in dialogue systems.
@@
    We propose an unsupervised framework, TG-DRG, designed to enhance the factual accuracy of responses without generation training. This framework adopts a textualised graph to generate responses built around reasoning-guided dialogue reformulation and dialogue sense knowledge selection. We validate TG-DRG against various comparable baselines; the experimental results show that TG-DRG can improve factuality noticeably.
@@
    We propose a supervised framework, GA-DRG. Different from TG-DRG, GA-DRG utilises a graph-based neural network to capture the node relationships, making it more efficient in improving factuality when generating dialogue responses. The experimental results show that GA-DRG outperforms the SOTA G-retriever in dialogue fact score with an improvement of 3.47% on OpendialKG and 3.12% on HybriDialogue.
Figure 2: The workflow of our proposed frameworks, TG-DRG and GA-DRG, begins with the dialogue. A CoT process first reformulates the dialogue. Based on the refined dialogue, the most valuable knowledge triples are then selected by the dialogue sense knowledge selection module. Starting from Step 4, the workflow diverges into two distinct paths for dialogue response generation: the blue solid arrows represent the TG-DRG path, while the purple dashed arrows indicate the GA-DRG path.
@@ 2.1 LLM Hallucinations and Mitigation Methods
LLMs are trending in NLP. However, hallucination exists in these LLMs widely, which is the phenomenon that language models generate coherent but fact-inconsistent text.
Several knowledge-based methods have been proposed to mitigate fact-conflicting hallucinations (Agrawal et al. 2023), which generally fall into three categories: knowledge-aware inference, training, and validation. For instance, Baek, Aji, and Saffari (2023) introduced KAPING, a prompt-based framework for QA that retrieves knowledge triples based on embedding similarity. Building upon retrieval techniques, Sen, Mavadia, and Saffari (2023) combined Knowledge Graph (KG) retrieval with language model reasoning to enhance performance on complex questions. Similarly, CRAG (Yan et al. 2024) employs an evaluator to assess generation quality and, if necessary, refines outputs via web search, showing significant improvements across various generation tasks. Another notable approach is Self-RAG (Asai et al. 2023), which retrieves relevant knowledge, generates an initial answer, and then refines it through self-critique, achieving strong results in multiple question answering benchmarks. In a related development, He et al. (2024) proposed G-Retriever, which integrates a graph-based encoder with the PCST algorithm to encode retrieved KG effectively, standing out among graph-enhanced knowledge retrieval methods.
However, applying QA-based approaches to multi-turn dialogue generation is limited by their lack of coreference modelling, which is essential for capturing discourse-level dependencies and ensuring factual consistency.
@@ 2.2 Evaluation of Factuality
Dialogue response generation not only needs to consider the dialogue context when generating responses but also needs to be evaluated based on the factuality and quality, rather than just the accuracy of answers of QA. Previous dialogue factuality evaluation mainly focuses on humans (Ni et al. 2023; Li et al. 2022; Yu et al. 2022), which is inefficient.
@@
Several QA datasets provide factual references in the form of entities; reference-based evaluation can then be based on entity matching. Other NLP datasets offer text references, so an alternative reference-based method involves matching the extracted entities between the generated text and these references (Nan et al. 2021). Instead, dialogue datasets provide reference responses. To evaluate the factual accuracy of the generated text, we can therefore compare the entities extracted from the generated response with those in the reference response, measured as the F1 score. However, relying solely on this can be inadequate, as the reference may not always provide explicit answers. It is crucial to consider additional metrics to ensure a comprehensive evaluation.
Reference-free methods can be categorised into uncertainty estimation (Farquhar et al. 2024) and external knowledge-based approaches, with the former often limited by its reliance on the generation model’s own confidence. The FactScore (Min et al. 2023), which falls into the latter category, is a metric designed to measure fact consistency in long-form text. The process begins with an LLM breaking down the text into fine-grained sentences called atomic facts. These atomic facts are then verified using both the LLM and external knowledge sources (Min et al. (2023) use Wikipedia titles). The FactScore is calculated based on the precision.
However, the FactScore is not directly suitable for our work for two reasons: First, it evaluates only the generated text without taking the dialogue history into account, which is crucial when assessing the dialogue factuality. Second, it does not account for the possibility that generated responses may lack external knowledge for verification.
@@ 3 Methodology
To address the limitations of prior works (He et al. 2024; Baek, Aji, and Saffari 2023), which often overlook complex pronoun resolution in dialogue settings and struggle to generate responses with high factual accuracy, we propose two novel frameworks: (1) Dialogue Response Generation via Textualised Graphs (TG-DRG) without training a generation model; and (2) Graph-Aware Dialogue Response Generation (GA-DRG), which leverages graph structures to better capture relationships between nodes. Both of them are built around a reasoning-guided coreference resolution and a dialogue sense knowledge selection. TG-DRG includes two variants based on the use of dialogue reformulation: TG-DRG-R incorporates a reformulation module, while TG-DRG-NR bypasses it. The overall workflow for our frameworks is illustrated in Figure 2.
In addition, to automatically evaluate factuality in dialogue systems, we propose the Dialogue Fact Score, a metric designed to assess factual accuracy specifically within dialogue contexts, thereby addressing the limitations of existing evaluation methods for open-domain dialogue factuality.
@@ 3.1 Task Formulation
We define the conversation history as a sequence of utterances $\mathcal{U}=\{\text{U}_{1},\text{U}_{2},\dots,\text{U}_{t-1}\}$, where $\text{U}_{i}$ denotes the utterance at the $i^{th}$ turn, either from the user or the system.
In our setup, external knowledge is modeled as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$ represents the set of nodes, which are entities, and $\mathcal{E}=\{e_{1},e_{2},\dots,e_{m}\}$ represents the set of edges, which are relations. The graph is constructed from an encyclopedia-based knowledge source, with each knowledge fact represented as a triple $(s,p,o)$, where $s,o\in\mathcal{V}$ denote the head and tail nodes, respectively, and $p\in\mathcal{E}$ denotes the edge (relation) connecting them.
Our task is to generate a factually correct response $\text{U}_{t}$ conditioned on the preceding utterances $\mathcal{U}$ and the knowledge graph $\mathcal{G}$.
3.2 Reasoning-Guided Dialogue and Knowledge Integration
We introduce a CoT-driven approach to dialogue reformulation, which systematically resolves coreferences step-by-step to produce high-quality dialogues. Additionally, we propose a dialogue sense knowledge selection method that constructs dialogue-relevant graphs to enhance factual accuracy. The reformulated dialogue enhances the effectiveness of knowledge selection and further mitigates hallucinations.
Reasoning-Guided Dialogue Reformulation
Zero-shot coreference resolution often struggles with resolving complex pronoun references in conversations due to the intricate reasoning involved. To address this challenge, we leverage the reasoning capabilities of LLMs to perform step-by-step resolution for each turn of the dialogue. This iterative approach not only simplifies the understanding of complex coreference relationships but also enhances the model’s overall coreference resolution ability. Specifically, we employ the CoT technique to guide the model through incremental reasoning, resulting in more accurate dialogue reformulation and improved factual consistency.
$\mathcal{U^{{}^{\prime}}}=f^{\text{ref}}_{\theta}(p_{\text{reform}},\mathcal{U}),$ (1)
where $p_{\text{reform}}$ is the prompt guiding the dialogue reformulation process, and $f^{\text{ref}}_{\theta}$ denotes an LLM that transforms the original dialogue $\mathcal{U}$ into its reformulated form $\mathcal{U}^{\prime}$.
Dialogue Sense Knowledge Selection
The external knowledge graph $\mathcal{G}$ contains noise and dialogue-irrelevant triples, making the LLM difficult to focus on the useful nodes when generating responses. To improve the quality, we propose a dialogue sense selection method relying on an LLM, which helps select the most dialogue-relevant triples and form a dialogue sense graph.
Specifically, we firstly encode the dialogue $\mathcal{U}$ and the triple $(s,p,o)$, respectively, into vector representations by the encoder $\psi$ and $\phi$. The function $f^{\text{sel}}_{\theta}$ denotes the LLM, which computes the score based on these representations. The formulation of selection probability is listed as follows:
$P(y=\text{rel}\mid\mathcal{U^{\prime}},(s,p,o))=\sigma\left(f^{\text{sel}}_{\theta}\left(\psi(\mathcal{U}^{(\cdot)}),\phi(s,p,o)\right)\right),$ (2)
where $\mathcal{U}^{(\cdot)}\in\{\mathcal{U},\mathcal{U}^{{}^{\prime}}\}$, rel denotes relevance. $\sigma$ is the activation function that maps the output to the probability.
We select the N triples based on the probability from the knowledge graph $\mathcal{G}$, and the process is defined as follows:
$\mathcal{G}_{\mathcal{U}^{(\cdot)}}=\text{Top}_{N}\left(\mathcal{E},\ \left\{P\left(\text{rel}\mid\mathcal{U}^{(\cdot)},\ (s,p,o)\right)\right\}_{(s,p,o)\in\mathcal{E}}\right),$ (3)
where $\mathcal{G}_{\mathcal{U}^{(\cdot)}}\in\{\mathcal{G}_{\mathcal{U}},\mathcal{G}_{\mathcal{U}^{{}^{\prime}}}\}$ is the dialogue sense graph constructed from the selected triples.
We select the LLaMA 3 8B model for dialogue reformulation and dialogue sense knowledge selection to achieve a balance between capability and efficiency. To enhance performance and leverage the strengths of advanced models, we fine-tune the LLaMA models using generated samples from GPT-4o, employing a combination of cross-entropy loss and Low-Rank Adaptation (LoRA) techniques (Hu et al. 2021).
3.3 Graph Knowledge-Driven Response Generation
The dialogue sense graph facilitates generating factual responses. We propose two graph-based generation methods.
Dialogue Response Generation via Textualised Graphs
Given the dialogue $\mathcal{U}^{(\cdot)}\in\{\mathcal{U},\mathcal{U}^{{}^{\prime}}\}$, as well as the dialogue sense graph $\mathcal{G}_{\mathcal{U}^{(\cdot)}}\in\{\mathcal{G}_{\mathcal{U}},\mathcal{G}_{\mathcal{U}^{{}^{\prime}}}\}$, the generation process is formulated as follows:
$\text{U}_{t}^{(\cdot)}=f^{\text{gen}}_{\theta}\big{(}p_{\text{gen}},\ \mathcal{U}^{(\cdot)},\ \mathcal{T}_{\mathcal{G}^{(\cdot)}}\big{)},$ (4)
where $p_{\text{gen}}$ denotes the prompt for generating the response. $f^{\text{gen}}_{\theta}$ is the LLM. We refer to the generation $\text{U}_{t}$ as TG-DRG-NR, and the generation $\text{U}_{t}^{{}^{\prime}}$ as TG-DRG-R. $\mathcal{T}_{\mathcal{G}^{(\cdot)}}$ denotes a textualised graph, formed of a set of text triples.
Graph Aware Dialogue Response Generation